OpenStudy (anonymous):

Find the exact value of the expression... sin (sin^-1(2/3) + cos^-1(1/3))

zepdrix (zepdrix):

AHHH sorry got busy for a few minutes there >.< Maybe I can help ya here.

zepdrix (zepdrix):

Confused on where to start?

zepdrix (zepdrix):

\[\Large\rm \sin\left[\color{orangered}{\sin^{-1}\left(\frac{2}{3}\right)}+\color{royalblue}{\cos^{-1}\left(\frac{1}{3}\right)}\right]\]So like....

zepdrix (zepdrix):

You know how the trig functions produce a ratio of sides? Example:\[\Large\rm \sin\theta=\frac{5}{7}\]We're taking the sine of some angle, and it gives us this ratio of opposite to hypotenuse.

zepdrix (zepdrix):

Inverse trig functions do the reverse. Example:\[\Large\rm \sin^{-1}\left(\frac{17}{4}\right)=\phi\]We're taking the inverse sine of some ratio of sides, and it spits out an angle.

zepdrix (zepdrix):

So that's how we want to deal with this problem. Think of the `orange` and `blue` and two different angles, and we're going to do some fancy stuff with them.

zepdrix (zepdrix):

So our problem really looks something like this:\[\Large\rm \sin\left[\color{orangered}{\theta}+\color{royalblue}{\phi}\right]\]Where,\[\Large\rm \color{orangered}{\theta=\sin^{-1}\left(\frac{2}{3}\right)}\]\[\Large\rm \color{royalblue}{\phi=\cos^{-1}\left(\frac{1}{3}\right)}\]

zepdrix (zepdrix):

We can rewrite these in terms of the normal trig functions,\[\Large\rm \color{orangered}{\sin \theta=\frac{2}{3}}\]\[\Large\rm \color{royalblue}{\cos \phi =\frac{1}{3}}\]

zepdrix (zepdrix):

And graph them on `separate` triangles using the side relations which are given.

zepdrix (zepdrix):

Do you see how we can use the 2/3 to label this triangle? Where you at Abbot, you around? :D